OpenStudy (anonymous):

The perimeter of the regular polygon shown is 14 ft. What is the side length of the polygon? A. 112 ft B. 3.5 ft C. 2.8 ft D. 1.75 ft

OpenStudy (anonymous):

Every regular polygon have equal magnitude of sides and angles. Since your polygon has 8 sides and a perimeter of 14 ft..let the side length of polygon =x ft. so perimeter =8*x ft or.. 8*x=14 x=14/8=1.75 ft So the edge length is (D) 1.75 ft

OpenStudy (anonymous):

do you not know how to do this or are you being lazy?

OpenStudy (anonymous):

I don't know